OpenStudy (anonymous):

Please answer I really need help!? Find the least common denominator of 3/x-1/6=7/12 and solve for x

OpenStudy (anonymous):

okay so your equation is \[\frac{ 3 }{ x }-\frac{ 1 }{ 6 }=\frac{ 7 }{ 12}\] so i would first start by adding 1/6 to both sides. 1/6 = 2/12, so you would end up with \[\frac{ 3 }{ x } = \frac{ 9 }{ 12 }\] well 9/12 = 3/4 so you have \[\frac{ 3 }{ x } = \frac{ 3 }{ 4 }\] so x must be 4

OpenStudy (anonymous):

And the least common denominator would be 9/12 or 3/4?

OpenStudy (anonymous):

to find the least common denominator of 3/x, 1/6 and 1/12, we must find common factor for all of them. well 6 factors into 12 nicely and 12x would be the smallest factor for x and 12 so 12x would be the least common denominator \[\frac{ 36 }{ 12x} - \frac{ 2x }{ 12x } = \frac{ 7x }{ 12x }\]

OpenStudy (anonymous):

this would be another way to solve the problem, because by finding the least common denominator, you can remove the fractions and end up with the equation 36 - 2x = 7x and if you solve for x again, it still equals 4

OpenStudy (anonymous):

Okay thank you so much! :)